can you help me solve this equation 7-3c=2(c+6)
and this please
a=3b+c making b the subject for the formula

thanks

Answer by stanbon(75887)

can you help me solve this equation
7-3c=2(c+6)
----
7-3c = 2c+12
5c = -5
c = -1
---------
and this please
a=3b+c making b the subject for the formula
3b = a-c
b = (a-c)/3
=====
Cheers,
Stan H.
============